# Learning Satellite Attitude Dynamics with Physics-Informed Normalising Flow

### v2

@@ 6 CONCLUSIONS
In this work, we introduced a novel framework for learning spacecraft attitude dynamics using a Real NVP neural network augmented with a self-attention mechanism and trained via a hybrid loss that combines data-driven supervision and physics-informed penalty. The relative weight of the two loss terms has been dynamically changed during training via the Lagrangian dual approach. Our analysis demonstrated that incorporating physical knowledge in the training as an inductive bias significantly enhances model performance, with improvements in terms of Mean Relative Error (MRE) ranging from 27.08% to 90.22% when forecasting 10 time steps. Embedding the learned models into an MPC framework further validated their practical utility, showing robust closed-loop control performance and resilience to observational noise. This evaluation was carried out by showing the trajectory distributions and by evaluating their average Mean Absolute Error (MAE), final error, and spread for both purely data-driven and physics-informed models. Additionally, the average absolute value of the Stability Performance Error (SPE) was computed to evaluate the stability of the controller under nominal conditions. Across all metrics, the physics-informed model consistently achieved the best results, demonstrating superior robustness and a SPE reduction of at least 16.67%. Although the best results were achieved with 10-step predictions, we believe that larger networks and longer horizons could yield even greater accuracy, though at increased computational cost. We believe that the proposed framework has promising generalization capabilities and could be used to address complex non-linear dynamics in future works, such as those involved in satellite berthing or docking with non-cooperative targets, or for the control of under-actuated spacecraft. These scenarios present greater modeling and control challenges due to factors like discontinuous dynamics, partial observability, and limited actuation, which align well with the strengths of combined data-driven and physics-informed learning.

### v3

@@ 6 CONCLUSIONS
In this work, we introduced a novel framework for learning spacecraft attitude dynamics using a Real NVP neural network augmented with a self-attention mechanism and trained via a hybrid loss that combines data-driven supervision and physics-informed penalty. The relative weight of the two loss terms has been dynamically changed during training via the Lagrangian dual approach. Our analysis demonstrated that incorporating physical knowledge in the training as an inductive bias significantly enhances model performance, with improvements in terms of Mean Relative Error (MRE) ranging from 27.08% to 90.22% when forecasting 10 time steps. Embedding the learned models into an MPC framework further validated their practical utility, showing robust closed-loop control performance and resilience to observational noise and RWs friction. This evaluation was carried out by showing the trajectory distributions of 300 Monte Carlo simulations and by evaluating their average steady-state errors, along with the settling time, for both purely data-driven and physics-informed models, and comparing their performances with traditional MPC solutions. Achieving an improvement in terms of settling time of about 62% compared to the best performing traditional MPC. Additionally, the Stability Performance Error (SPE) was computed to evaluate the stability of the proposed controller. Across all metrics, the physics-informed model consistently achieved the best results compared to the data-driven approach, demonstrating superior robustness. Although the best results were achieved with 10-step predictions, we believe that larger networks and longer horizons could yield even greater accuracy, though at increased computational cost. We believe that the proposed framework has promising generalization capabilities and could be used to address complex non-linear dynamics in future works, such as those involved in satellite berthing or docking with non-cooperative targets, or for the control of under-actuated spacecraft. These scenarios present greater modeling and control challenges due to factors like discontinuous dynamics, partial observability, and limited actuation, which align well with the strengths of combined data-driven and physics-informed learning. Finally, we envision the integration of traditional and AI-based approaches to leverage the strengths of both, enabling faster and more accurate attitude control, even under non-nominal conditions. In the context of this work, a potential robust solution may see the integration of the MPC with physics-informed dynamics for the highly dynamic maneuvers and steady-state dynamics when close to the target attitude.